Mission Operations Days: 2001/286 to 2001/292
October 13, 2001 00:00:00 GMT to October 19, 2001 00:00:00 GMT
NOTE: A new Table #5: Ideal Blackbody Output, was loaded into the instrument on Friday, October 12, 2001 (2001/285) at approximately 13:45:00 GMT. The following changes were made:
Lowered values for Bands 5 to 7 by 200 (decimal) counts.
Increased values for Bands 20, 22 to 25, 27 and 28 by 100 counts
If the new value would have been larger than 4095, then the value was set to 4095. This setting occurred for higher temperature values of B24, B25, B27, and B28.
Data Loss on DOY 292
See “Non-MODIS Significant Events” below for details.
Terra Spacecraft and MODIS Instrument Status:
|Terra (AM-1) is in Normal Ops Mode|
MODIS is in Science Mode - A Side
|Blackbody||A On (270 K); B Off||Nominal|
|Calibration Electronics||A On; B Off||Nominal|
|Control Processor||A On; B Off||Nominal|
|Nadir Door||Unlatched, Open||Nominal|
|Space View Door||Unlatched, Open||Nominal|
|Solar Diffuser Door||Unlatched, Closed||Nominal|
|FDDI Formatter||A On; B Off||Nominal|
|FDDI Port||A On; B Off||Nominal|
|FIFO Memory||#1 & #2 On; #3 & #4 Off||Nominal|
|Format Processor||A On; B Off||Nominal|
|Power Supplies||#1 On; #2 Off||Nominal|
|PV VIS FPA||A On; B Off||Nominal|
|PV NIR FPA||A On; B Off||Nominal|
|PV SMIR FPA||A On; B Off||Nominal|
|PV LWIR FPA||A On; B Off||Nominal|
|PC LWIR FPA||A On; B Off||Nominal|
|RC Outgas Heaters||All Off (CS, IS, OS)||Nominal|
|RC LWIR FPA Heater||On (83 K)||Nominal|
|RC SMIR FPA Heater||Off||Nominal|
|Scan Assembly||A On; B Off||Nominal|
|SDSM||A Off; B Off||Nominal|
|SRCA||A Off; B Off||Nominal|
|PS1 Survival Heater||Enabled||Nominal|
|PS2 Survival Heater||Enabled||Nominal|
|Timing Generator||A Off; B On||Nominal|
|Flight Software||Rev BD + 3 set of patches||Nominal|
|Inhibit Ids Set||None||Nominal|
|TMON 66 (SRCA A)||Disabled||Nominal|
|TMON 67 (SRCA B)||Disabled||Nominal|
|TMON 68 (FIFO)||Enabled||Nominal|
|TMON 69 (EPS Current)||Disabled||Nominal|
This Week's Completed MODIS Activities:
Saturday, October 13, 2001
Sunday, October 14, 2001
Monday, October 15, 2001
Tuesday, October 16, 2001
Wednesday, October 17, 2001
2001/290 16:11:37-16:23:05 ATC OA15: SD/SDSM Open
2001/290 17:50:30-18:01:58 ATC OA16: SD/SDSM Screened
2001/290 21:00:00 ATC OA26: BB Temperature Stepping (BB to 270K)
Thursday, October 18, 2001
2001/291 15:00:00 ATC OA26: BB Temperature Stepping (BB to 280K)
2001/291 19:00:00 ATC OA26: BB Temperature Stepping (BB to 285K)
2001/291 19:07:39 Real-time Memory dump of formatter error event addresses
2001/291 19:08:42 Real-time Memory dump of formatter error event counts
2001/291 22:00:00 ATC OA26: BB Temperature Stepping (BB to 290K)
Friday, October 19, 2001
2001/292 02:00:00 ATC OA26: BB Temperature Stepping (BB to 295K)
2001/292 02:00:30 ATC OA26: BB Temperature Stepping (BB Duty Cycle to FULL)
2001/292 04:00:00 ATC OA26: BB Temperature Stepping (BB to 300K)
2001/292 06:00:00 ATC OA26: BB Temperature Stepping (BB to 315K)
2001/292 09:00:00 ATC OA26: BB Temperature Stepping (BB to 270K)
2001/292 09:00:30 ATC OA26: BB Temperature Stepping (BB Duty Cycle to THIRD)
This Week's Scheduled MODIS Activities Not Completed:
Upcoming MODIS Events:
Saturday, October 20, 2001
293/05:00:00 ATC OA26: BB Temperature Stepping (BB to 290K)
Two SD/SDSM calibrations will be scheduled during the upcoming week.
The next Drag Makeup has officially moved from 2001/298 to 2001/303 (October 30, 2001), around 1900z. An official time has yet to be published.
The next Lunar Calibration roll maneuver is scheduled for day 2001/309 (November 5, 2001) with data collection from 18:15:11 to 18:21:41. The roll angle will be -11.8237 degrees.
The topics of Terra's orbit inclination adjust maneuvers and eventual equator crossing time has resurfaced. Terra's equator crossing mean local time (MLT) requirement is 10:30 +/- 00:15. The current crossing time is about 10:35 and decreasing by 1-2 seconds per day. The orbit management talks have begun because, under current conditions, Terra will approach the SAC-C spacecraft very closely in April of 2002. Further details will follow as a finalized plan emerges.
Formatter Error Events:
The FR error counter (MOD_SS_CP_STATUS_08) has “rolled over” 30 times since the A-Side restart following the PS-2 Anomaly, and 28 times since the latest string of errors started on August 15. This is a 16-bit signed counter, therefore it “rolls over” every 32,768 counts.
Following last week's brief respite during the SRCA calibrations, the formatter error rate returned to its previous level at about 2,000-2,500 errors per hour and continued on its slowly increasing trend.
As part of a blackbody calibration currently underway, the BB temp is lowered from 290K to 270K. This has caused the FR_A_ENGINE temperature to drop slightly (to around -0.5 C) and the error rate increased in turn to about 4,000 per hour, causing the event counter to rollover in only 8 hours. The BB then began its scheduled heating in various temperature steps from 280K to 315K. During that period, the error rate dropped to below 1000 errors per hour when the BB was at the upper temperature range (and the associated heating brought the FR_A_ENGINE to around +1.0 C). The BB then began cooling to 270K, and as a result, the FR_A_ENGINE temperature decreased and the error rate once again increased. We should expect an increased error rate until after 05:00 GMT Saturday when the BB is returned to the nominal 290K operating temperature.
Two error tracking spreadsheets have been updated and can be retrieved from:
An MS Excel spreadsheet called, AM1_FR_error_rollovers_101801.xls is the most recent version, with the latest data available. A second spreadsheet, MOD_ME_Orbit_temps_2001_195-285.xls contains 90 days of orbital average temperatures for a wide variety of main electronics module mnemonics. Only one plot (of the FR A/B temperatures) has been included, but the data are available for other such plots.
General Instrument Comments:
MODIS is in Science Mode on the A-side with the SVD and NAD open, operating with the Formatter Resolution Flight Software Patches (FRFSP).
Note: SD/SDSM calibrations were planned to be reduced to once per two weeks. This plan has been revised; SD/SDSM calibrations will continue to be scheduled once each week until the end of November.
MODIS Telemetry Trends:
Instrument temperatures across the board are rising slightly as of late, although the orbit beta angle has not yet reached its minimum.
Non-MODIS Significant Events:
The Network Communications Center (NCC) has been rejecting and/or canceling Terra TDRSS contact times (likely to support DOD satellites). This has resulted in at least one incidence of data loss.
On DOY 292, an upcoming (post 20:00:00 GMT) TDRSS contact was switched on short notice (for the reasons described above). The electronic tables sent to EDOS were not updated in time to include this information. As a result, SSR partition management was incapable of correctly compensating for the necessary data playbacks. The MODIS (and other) SSR partitions will fill and further Science Data will not be recorded for an unspecified duration of time. Expected data loss is on the order of one orbit's worth. Specific details will be forwarded when they become available from the FOT.
Limited Life Item Status:
SRCA 10W Lamp #1: 192.2 of 500 hours
SRCA 10W Lamp #2: 146.6 of 500 hours
SRCA 10W Lamp #3: 155.8 of 500 hours
SRCA 10W Lamp #4: 61.5 of 500 hours
SRCA 1W Lamp #1: 558.3 of 4000 hours
SRCA 1W Lamp #2: 276.3 of 4000 hours
Solar Diffuser Door: 1733 of 3022 Movements
Nadir Aperture Door: 532 of 1316 Movements
Space View Door: 437 of 1316 Movements